find the equation for the line that is parallel to the line wih the equation (y+2)=1/2 (x-4) passing through the point (1,3)
Answer by Maths68(1474)


Standard Form of Equation of the line:
y=mx+b
Given
(y+2)=1/2 (x-4)
rearrage the above equation according to the standard form
y+2=(1/2)x-4(1/2)
y+2=(1/2)x-2
y=(1/2)x-2-2
y=(1/2)x-4
Compare above equation with the standard form equation
m=1/2 and b=-4
Since lines are parallel their slopes will be equal,therefore slope of the required line will be (1/2)
Now we have a point(1,3) and slope (1/2)of the line we can easily find required lines by putting these values in the equation of the straight line poin-slope form.
m=(y2-y1)/(x2-x1)
1/2=(y-3))/(x-1)
1*(x-1)=2(y-3)
x-1=2y-6
x-1+6=2y
x+5=2y
2y=x+5
2y/2=(x+5)/2
y=(1/2)x+5/2
Above equation is the Equation of Staright line in the Standard Form.